OpenStudy (anonymous):

Gordon's club wants to donate 42 cans of food to the community food pantry. The club has collected 18 cans of soup and 15 cans of vegetables. How many more cans of food does the club need to collect? A. Add 18 and 15. Subtract this sum from 42. Gordon's club has to collect 9 more cans. B. Add 18 and 15. Add this sum to 42. Gordon's club has to collect 75 more cans. C. Subtract 15 from 18. Add this difference to 42. Gordon's club has to collect 45 more cans.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

If you you need $42 and I give you $18 and $15, how much more money do you need @cute-girl ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

its b?

OpenStudy (anonymous):

do you know

OpenStudy (skullpatrol):

$$\Huge 42-18-15$$$$=42+(-18)+(-15)$$

OpenStudy (skullpatrol):

$$\Huge a-b=a+(-b)$$

OpenStudy (skullpatrol):

To subtract a number, add the opposite of that number.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok